# The thermodynamic structure and large-scale structure filament in MACS J0717.5+3745

### v2

@@ 3.2 XMM-Newton Instrumental Background
The FWC data were modelled in the 0.3 keV to 10 keV band for the MOS1 and MOS2 detectors and in the 0.4 keV to 10 keV range for pn. The continuum emission is fitted by a broken power-law, while all elemental lines were modelled using Gaussians described in Tab. 3. Because these HP components are noise contributions directly from the detectors and are unrelated to the mirrors, these instrumental background models are not convolved by the Ancillary Response File (ARF).
@@
To further check for any residual soft proton contamination, we compared the area-corrected count rates between the ’inside’ and ’outside’ regions of the Field of View (FOV) for each detector in each observation, shown in Tab. 4 (De Luca & Molendi, 2004). All of the diagnostic values are under the lowest soft proton threshold of 1.15, indicating that none of the event lists are additionally contaminated by soft protons. As such, in this paper, the soft proton parameters were removed from the background model, which reduced the overall complexity. All XMM-Newton spectral fits reported in this paper were performed in the 0.45 to 7.0 keV range for MOS and 0.3 to 7.0 keV for pn.
@@ 3.3 The Astrophysical X-ray Background
The Cosmic X-ray Background (CXB) includes the unresolved point sources, distant Active Galactic Nuclei (AGN) typically modelled with a photoelectrically absorbed power-law with slope frozen at 1.41 and a free normalization (Hickox & Markevitch, 2006). The Galactic foreground sources can be described as a combination of several plasmas in collisional ionisation equilibrium (apec). They include the Local Hot Bubble (LHB), the Galactic Halo (GH), and recently another component, interpreted as a super-virialized (SV) component of our Galactic Halo, sometimes referred to as the ‘Galactic Corona’. The LHB is an approximately 0.1 keV plasma originating from a local supernova remnant that we are currently residing in. It can be effectively modelled using the ROSAT All Sky Survey (RASS) data with the support of the soft X-ray bands from Chandra and XMM-Newton. Additionally, our Galactic Halo and its supervirial component are described by photoelectrically absorbed (tbabs, wilms), thermal APEC plasma models (Wilms et al., 2000). For all of these local, foreground CXB components, the redshift was frozen to 0, the abundances frozen to solar, with the temperatures and normalizations free to fit. Unless otherwise stated, the 1.41 power law photon index was kept frozen. Meanwhile, the neutral hydrogen column density for the CXB and in all subsequent fits was frozen to $N_{\text{H}}$ = $8.36\times 10^{20}$ cm-2, corresponding to the sum of the atomic and molecular hydrogen column densities along the line of sight [^1] https://www.swift.ac.uk/analysis/nhtot/ (Kalberla et al., 2005; Willingale et al., 2013).
@@
The top part of Tab. 5 shows the results of the different CXB components, with and without the inclusion of the ROSAT All Sky Survey data, using both a background subtracted and background modelled fits. Without the ROSAT data, the LHB temperature was fixed to 0.09 keV and a free normalization. Using all available data and the background models, the best-fit results using all 13 spectra gave a LHB temperature of 0.108±$\pm 0.02$ keV, GH temperature of 0.156±$\pm 0.002$ keV, and SV temperature of $0.697\pm 0.011$ keV, with a SV normalization around 10% that of the GH, consistent with other studies.
In all of the spectral analysis in this paper, spectral data was never combined, but instead, all data was joint fit simultaneously. For each of the spectral data groups, a 20% systematic error was applied to the data in Sherpa/Xspec to account for any cross-calibration related discrepancies in measurements between the instruments.
@@ 4.1 Thermodynamical Maps
Once the CXB parameters are constrained, it is possible to more accurately constrain the cluster emission and, consequently, to create thermodynamical maps and trend-averaged maps of MACS J0717.5+3745 to better study and visualize the ICM physics and internal structures. Specifically, we used a modified version of the contour binning algorithm from Sanders (2006), which we first translated into Python from C++, and have released as pycontbin [^3] https://github.com/jpbreuer/contbin_python. This contour binning algorithm creates regions with a pre-selected S/N ratio while conserving the general contours of the data by grouping neighbouring pixels of similar surface brightness and yields statistically independent regions. The contour binning maps were derived from an exposure corrected XMM-Newton mosaic image, scaled to have the approximate number of counts per pixel as the sum of observations. The smallest bin in this map corresponds to a circle of around 10.6 arcseconds radius, around the size of the Point Spread Function (PSF) of XMM-Newton. Like the CXB region, all contour bin map regions were converted into polygons and were extracted with both Chandra and XMM-Newton, with the idea that the Chandra resolution could be exploited in the joint fitting procedure as a correction to the fit results in the cases where the complex polygon region geometries occasionally fall under the size of the XMM-Newton psf. This resulted in a joint fit between 25 different spectra for each spatial bin, 13 from the CXB region, and 12 spectra from each spatial bin.
We again use the model for a plasma in collisional ionisation equilibrium (apec) with photoelectric absorption (tbabs), but now with fixed values for the metallicity at 0.3 Solar and redshift at $0.5458$. We additionally scale each region by its area in arcminute2 using a model constant, so that we can simultaneously fit the large cxb region with each cluster spectra. The CXB parameters are first frozen to their best-fit values before fits are performed to constrain the cluster temperatures kT and normalizations norm. Finally, the ICM spectra abundances are thawed to constrain the metallicity parameter for each region. This fitting procedure is repeated for every bin and for every instrument independently before being performed on the combined data from all observatories and again for the combined data from all available instruments. Unless otherwise mentioned, this spectral fitting methodology applies to any other spectral analysis considered for the rest of the paper.
@@
[…]
@@
Due to the nature of merging systems, projection effects may bias the results due to potential asymmetries along the projected line of sight. Moreover, because this cluster is a highly-disturbed, non-spherically symmetric, merging cluster with possibly more than 4 distinct sub-cluster populations in different merging states, it is difficult to easily approximate the integration volume due to uncertainties in the projected geometry. As a simplification, we assume a flat, pancake-like approximation for the line of sight depth of 1 Mpc for each map pixel in each bin. Introducing more complex geometric assumptions, eg. several beta models, or even a spherically symmetric model, would introduce more uncertainties to the maps, and would simultaneously increase the difficulty in interpreting the features observed in the maps (Veronesi et al., 2024). In exchange for some control over the geometric uncertainty, the consequence of the flat 1 Mpc approximation is that electron density values near the cluster center might be underestimated.
@@ 4.2 Azimuthal Trend-divided Maps
Figure 5: Trend-divided maps of temperature (top left), density (top right), pressure (bottom left), and entropy (bottom right).
@@ 4.3 Spectroscopic Disentanglement of the Filament Signal
The spectroscopic fitting of the filamentary region was meticulously performed in several steps. We keep the standard assumption that it is a faint signal contaminated by instrumental backgrounds, as well as both astrophysical foreground and background signals; however, we extend the astrophysical description of the background to also include the local atmospheres of nearby systems. The filament lies between MACS J0717.5+3745 and a group to the southeast, inside their respective $R_{200}$ emission areas, so to properly constrain the temperature and density of the filament, it is first necessary to disentangle the different contributions to the complex signal.
To constrain the cluster and group-related contributions to the filament signal, large arc regions are extracted, approximately at the same radial distance at which the filament structure is displaced from the cluster and group centers, inside, but close to R200. These regions are then independently fit using a similar method described in Sec. 4.1, using CXB parameter constraints as determined from the optimal CXB fit, and then applying these posteriors as priors to the subsequent fits. The fit results of the CXB + Cluster and CXB + Group regions are shown in the center of Tab. 5.
As we do not know the exact multi-phase composition of the filamentary gas, we try to statistically disentangle the multi-phase nature of the region using Bayesian X-ray Analysis and model selection using the AIC and BIC criteria described in Appendix Sec. A. Tab. 6 shows the AIC and BIC values when comparing a combination of different spectral models that could effectively describe the filament region. To fit all of the various models, we used only the CXB and filament spectra to fit the temperature and normalization of the filament data while all other parameters were frozen to their best fits. In these data were 3116 data points, and approximately 3114 degrees of freedom, so the AIC and BIC values are fairly close to each other.
@@
Based on the different criteria, all of the various apec models were close enough in their likelihoods with each other that they were not outright rejected and theoretically could each be effectively used to describe the data. That being said, the most complicated of the models, which contained the CXB + Cluster + Group + Filament, ended up with the best fit results, despite being punished for its extra complexity. Interesting to note is that the ‘second-best’ model according to the AIC is the CXB + Cluster + Filament, and not the expected CXB + Group + Filament model. This could be due to a bias in the region selection area, where we might have included more cluster emission than group emission, that the group atmosphere is not being effectively stripped by the filament accretion, that the filament region could be representative of a more simple ‘mixing’ zone between the two atmospheres, and there is a stronger cluster signal due to the difference in the size scales of the respective atmospheres, or due to other projection effects.
Figs. 6,7,8 show the source and background spectral data from Chandra, XMM-Newton MOS1 and MOS2, and XMM-Newton pn of the filament region. Besides the full astrophysical and instrumental background models, each plot also shows the individual contributions to the total model from each astrophysical temperature component, along with a residual plot for the spectral data and the full model, and a second residual plot for the instrumental background model. The counts for each instrument are averaged across each energy bin, and then binned for better visualization. As can be seen in the Figures, the contribution from each temperature component was successfully constrained by the fitting.
@@ 4.4 The Sunyaev-Zeldovich Effect
[…]
@@
where $x=\frac{h\nu}{k_{B}T_{\rm{CMB}}}\approx\nu/56.8\rm{GHz}$ is the dimensionless frequency, $h$ is the Planck constant, $\nu$ is the observed frequency, $k_{B}$ is the Boltzmann constant, and $T_{\rm{CMB}}=2.7255\rm{K}$ is the CMB temperature. The $\delta_{\rm{tSZ}}(x,T_{e})$ and $\delta_{\rm{kSZ}}(x,v_{z},T_{e})$ terms correspond to the relativistic corrections for the tSZ and kSZ effects, and are dependent on the observing frequency and $T_{e}$, and in the kSZ case, also the line of sight velocity. The relativistic corrections for the tSZ are computed using the work of Itoh & Nozawa (2003), while the kSZ corrections are computed using the analytical formula in Nozawa et al. (2006).
@@ 4.5 Correction of kinetic Sunyaev-Zeldovich via ICM Physical Modeling
Because MUSTANG2 is dominated by the tSZ and limited to a single frequency $\sim 90\rm{GHz}$, it does not have much constraining power for the kSZ signal. To build a kSZ model for MUSTANG2 to make the correction, we use the same model parameters as ‘F2’ constrained in Adam et al. (2017b), which used the NIKA data and extra information from X-ray imaging to break degeneracies between $v_{z}$ and $\tau_{e}$. We start with the fact that previous strong lensing observations in Limousin et al. (2016) have shown that there are at least four subclusters present in MACS J0717.5+3745, listed in Tab. 7, and shown as circles in Fig. 9. Due to the linear dependence of $n_{e}$ in the SZ data, the 3D gas distribution is built as a sum of each independent subcluster, where each subcluster is modeled with a spherically symmetric $\beta$-model,
@@
and $n_{e0}$ is the central electron density, $r_{c}$ is the core radius, and $\beta$ is the slope of the radial profile (Cavaliere & Fusco-Femiano, 1978). Each subcluster additionally is assumed to have isothermal gas temperatures, determined using 30 arcsecond regions from XMM-Newton in Adam et al. (2017b), and constant bulk velocities, all given in Tab. 7. This gas distribution is then analytically integrated along the line of sight for each subcluster as,
@@
[…]
where $T_{x}$ and $v_{z}$ are scalar quantities for each subcluster.
Fig. 9 shows a comparison plot between a zoomed in view of the X-ray derived pressure and the trend-divided pressure maps (top), followed by Sunyaev-Zeldovich MUSTANG2 data, and a kSZ and rSZ corrected version of the data. The white box corresponds to the FOV of the MUSTANG2 data, and is the same as the thermodynamical maps shown in Fig. 4 and Fig. 5, while the new smaller green box shows the zoomed in region in the two top plots. The white contours are the radio images in van Weeren et al. (2017).
There is a relatively close agreement between the X-ray pressure peak and the SZ peak, with an offset corresponding to around 0.25 arcminutes, or around 100 kpc at this redshift, but this estimate is biased due to the resolution of the MUSTANG2 data and the size and shape of the corresponding bin from the X-ray pressure map. As can be seen from the trend-divided pressure map, there is also a clear offset between the radio relic emission and the X-ray pressure peak, likely due to the nature of the integrated X-ray emission and projection effects mentioned in van Weeren et al. (2017). We note that the kinetic and relativistic correction appears to have a larger than expected effect on the MUSTANG2 data, this could be attributed to the large uncertainties associated with the model.
As mentioned previously, the present work does not explore pressure fluctuations (see Khatri & Gaspari, 2016; Romero et al., 2023b, 2024), but we note that deeper SZ observations with, for example, MUSTANG2, could improve upon the constraints for MACS J0717.5+3745 presented in Adam et al. (2024). Future MUSTANG2 observations could also provide a low frequency ballast point for constraining cluster temperatures via the relativistic correction to the thermal SZ, which, along with the known kinetic SZ shift, may explain the apparent discrepancy between X-ray and SZ results.
Figure 9: Comparison plots between the X-ray pressure map (top left), the trend-divided X-ray pressure map (top right), followed by Sunyaev-Zeldovich MUSTANG2 data, displayed here as a SNR image (bottom left), next to a kinetic and relativistic SZ corrected version of the data, shown here in KelvinCMB units (bottom right). The white contours are from the radio images seen in van Weeren et al. (2017), and the green contours corresponds to the MUSTANG2 data in the bottom left. The circles represent the strong lensing subcluster centers given in Limousin et al. (2016) which were used to reproduce the kSZ+rSZ model used in Adam et al. (2017b). The full FOV of the MUSTANG2 data can be seen in Appendix Sec. C.

### v3

@@ 3.2 XMM-Newton Instrumental Background
The FWC data were modelled in the 0.3 keV to 10 keV band for the MOS1 and MOS2 detectors and in the 0.4 keV to 10 keV range for pn. The continuum emission is fitted by a broken power-law, while all elemental lines were modelled using Gaussians described in Table 3. Because these HP components are noise contributions directly from the detectors and are unrelated to the mirrors, these instrumental background models are not convolved by the Ancillary Response File (ARF).
@@
To further check for any residual soft proton contamination, we compared the area-corrected count rates between the ’inside’ and ’outside’ regions of the Field of View (FOV) for each detector in each observation, shown in Table 4 (De Luca & Molendi, 2004). All of the diagnostic values are under the lowest soft proton threshold of 1.15, indicating that none of the event lists are additionally contaminated by soft protons. As such, in this paper, the soft proton parameters were removed from the background model, which reduced the overall complexity. All XMM-Newton spectral fits reported in this paper were performed in the 0.45 to 7.0 keV range for MOS and 0.3 to 7.0 keV for pn.
@@ 3.3 The Astrophysical X-ray Background
The Cosmic X-ray Background (CXB) includes the unresolved point sources, distant Active Galactic Nuclei (AGN), typically modelled with a photoelectrically absorbed power-law with slope frozen at 1.41 and a free normalization (Hickox & Markevitch, 2006). The Galactic foreground sources can be described as a combination of several plasmas in collisional ionisation equilibrium (apec). They include the Local Hot Bubble (LHB), the Galactic Halo (GH), and recently another component, interpreted as a super-virialized (SV) component of our Galactic Halo, sometimes referred to as the ‘Galactic Corona’. The LHB is an approximately 0.1 keV plasma originating from a local supernova remnant that we are currently residing in. It can be effectively modelled using the ROSAT All Sky Survey (RASS) data with the support of the soft X-ray bands from Chandra and XMM-Newton. Additionally, our Galactic Halo and its supervirial component are described by photoelectrically absorbed (tbabs, wilms), thermal APEC plasma models (Wilms et al., 2000). For all of these local, foreground CXB components, the redshift was frozen to 0, the abundances frozen to solar, with the temperatures and normalizations free to fit. Unless otherwise stated, the 1.41 power law photon index was kept frozen. Meanwhile, the neutral hydrogen column density for the CXB and in all subsequent fits was frozen to $N_{\text{H}}$ = $8.36\times 10^{20}$ cm-2, corresponding to the sum of the atomic and molecular hydrogen column densities along the line of sight [^1] https://www.swift.ac.uk/analysis/nhtot/ (Kalberla et al., 2005; Willingale et al., 2013).
@@
[…]
In all of the spectral analysis in this paper, spectral data was never combined, but instead, all data was fit simultaneously. For each of the spectral data groups, a 20% systematic error was applied in Sherpa/Xspec to account for any cross-calibration related discrepancies in measurements between the instruments. A more detailed discussion regarding the cross-calibration can be found in Appendix B.
Once the CXB parameters are constrained, it is possible to more accurately constrain the cluster emission and, consequently, to create thermodynamical maps and trend-averaged maps of MACS J0717.5+3745 to better study and visualize the ICM physics and internal structures. Specifically, we used a modified version of the contour binning algorithm from Sanders (2006), which we first translated into Python from C++, and have released as pycontbin [^3] https://github.com/jpbreuer/contbin_python. This contour binning algorithm creates regions with a pre-selected S/N ratio while conserving the general contours of the data by grouping neighbouring pixels of similar surface brightness and yields statistically independent regions. An added benefit of this modified contour binning code is that it can additionally take a static value for a Point Spread Function (PSF) as a constraint and merge map bins that are smaller than the given value, ensuring bins larger than the instrument PSF. The contour binning maps were derived from an exposure corrected XMM-Newton mosaic image, scaled to have the approximate number of counts per pixel as the sum of observations. The smallest bin in this map corresponds to a circle of around 10.6 arcseconds radius, around the size of the PSF of XMM-Newton. Like the CXB region, all contour bin map regions were converted into polygons and were extracted with both Chandra and XMM-Newton, with the idea that the Chandra resolution could be exploited in the joint fitting procedure as a correction to the fit results in the cases where the complex polygon region geometries occasionally fall under the size of the XMM-Newton psf. This resulted in a joint fit between 25 different spectra for each spatial bin, 13 from the CXB region, and 12 spectra from each spatial bin.
We again use the model for a plasma in collisional ionisation equilibrium (apec) with photoelectric absorption (tbabs), but now with fixed values for the metallicity at 0.3 Solar and redshift at $0.5458$. We additionally scale each region by its area in arcminute2 using a model constant, so that we can simultaneously fit the large cxb region with each cluster spectra. The CXB parameters are first frozen to their best-fit values before fits are performed to constrain the cluster temperatures kT and normalizations norm. Finally, the ICM metallicity is thawed to constrain its value for each region. This fitting procedure is repeated for every bin and for every instrument independently before being performed on the combined data from all observatories and again for the combined data from all available instruments. Unless otherwise mentioned, this spectral fitting methodology applies to any other spectral analysis considered for the rest of the paper.
@@ 4.1 Thermodynamical Maps
$\textit{norm}=\frac{10^{-14}}{4\pi(d_{\textsc{a}}(z+1))^{2}}\int n_{\textrm{e}}n_{\textrm{p}}dV\,$, (1)
@@
Given the inherent unknowns and uncertainties associated with the true three-dimensional geometry of MACS J0717.5+3745, we explore the systematic effects resulting from the simplifying assumption of a constant 1 Mpc line of sight depth in our original thermodynamic maps. To quantify potential biases introduced by this approximation, we implement a geometrical correction derived from a model of the cluster in a 3D box. Specifically, we generated an effective line-of-sight length map using 5 different 3D spherical beta models parameterized according to the lensing-detected subclusters listed in Table 8, with an additional beta model added to include the group at the end of the filament. The best fit results are given in Table 5. The emissivity distribution was then integrated from the center of the 3D simulated cube until 50% of the total flux along the line-of-sight was reached, shown in Fig. 5.
Table 5: Best‐fit parameters for five spherical beta‐models determined from Chandra data. The central positions for subclusters A-D are the same shown in Table 8, and are determined form the strong lensing centers in Limousin et al. (2016), while the center of subcluster E is fixed to the center of the group region.
Sub. Amp. $r_{c}$ [kpc] $\beta$ RA Dec
A 0.577 785 1.855 07:17:25.0 +37:45:54.6
B 1.483 279 0.896 07:17:31.3 +37:45:30.3
C 0.851 166 0.624 07:17:35.8 +37:45:01.0
D 0.977 145 0.577 07:17:33.0 +37:44:15.0
E 0.549 69 0.495 07:17:53.3 +37:42:11.8
Figure 5: Map of the line of sight effective length derived from the 3D numerical simulation. The five subcluster centers are shown as white circles, and correspond to the different beta models shown in Table 5.
Furthermore, we quantify the effect of the geometric correction used in this paper with additional maps of the fractional change and associated sigma change, using the propagated uncertainties derived from spectral fitting for each region, which can be seen in Appendix D Fig. 16 and Fig. 17.
Figure 6: Maps of temperature (top left), density (top right), pressure (bottom left), and entropy (bottom right) with corresponding units of keV, cm-3, keVcm-3, and keVcm2. The temperature and electron density maps were created using binning with a S/N of 120 (14400 counts) and 100 (10000 counts), respectively, while the pressure and entropy maps were produced using the results from the top two S/N maps. Density maps were derived from the normalization fits assuming an effective length determined from the best fit results of 5 beta models, shown in Table 5. Spectral fits were made using data from both Chandra and XMM-Newton for the temperature, and only Chandra for the density, each using the CXB best fit results from ROSAT, Chandra, and XMM-Newton. Fits for these maps were performed using the most appropriate background model for each region according to the AIC.
@@ 4.2 Azimuthal Trend-divided Maps
Figure 7: Geometry-corrected, trend-divided maps of temperature (top left), density (top right), pressure (bottom left), and entropy (bottom right).
@@ 4.3 Spectroscopic Disentanglement of the Filament Signal
The spectroscopic fitting of the filamentary region was performed in several steps. We keep the standard assumption that it is a faint signal contaminated by instrumental backgrounds, as well as both astrophysical foreground and background signals; however, we extend the astrophysical description of the background to also include the local atmospheres of nearby systems. The filament lies between MACS J0717.5+3745 and a group to the southeast, inside their respective $R_{200}$ emission areas, so to properly constrain the temperature and density of the filament, it is first necessary to disentangle the different contributions to the complex signal.
To constrain the cluster and group-related contributions to the filament signal, large arc regions are extracted, approximately at the same radial distance at which the filament structure is displaced from the cluster and group centers, inside, but close to R200. These regions are then independently fit using a similar method described in Sec. 4.1, using CXB parameter constraints as determined from the optimal CXB fit, and then applying these posteriors as priors to the subsequent fits. The fit results of the CXB + Cluster and CXB + Group regions are shown in the center of Table 6.
As we do not know the exact multi-phase composition of the filamentary gas, we try to statistically disentangle the multi-phase nature of the region using Bayesian X-ray Analysis and model selection using the AIC and BIC criteria described in Appendix Sec. A. Table 7 shows the AIC and BIC values when comparing a combination of different spectral models that could effectively describe the filament region. To fit all of the various models, we used only the CXB and filament spectra to fit the temperature and normalization of the filament data while all other parameters were frozen to their best fits. In these data were 3116 data points, and approximately 3114 degrees of freedom, so the AIC and BIC values are fairly close to each other.
@@
Based on the different criteria, all of the various apec models were close enough in their likelihoods with each other that they were not outright rejected and theoretically could each be effectively used to describe the data. That being said, the most complicated of the models, which contained the CXB + Cluster + Group + Filament, ended up with the best fit results, despite being punished for its extra complexity. Interesting to note is that the ‘second-best’ model according to the AIC is the CXB + Cluster + Filament, and not the expected CXB + Group + Filament model.
Figs. 8,9,10 show the source and background spectral data from Chandra, XMM-Newton MOS1 and MOS2, and XMM-Newton pn of the filament region. Besides the full astrophysical and instrumental background models, each plot also shows the individual contributions to the total model from each astrophysical temperature component, along with a residual plot for the spectral data and the full model, and a second residual plot for the instrumental background model. The counts for each instrument are averaged across each energy bin, and then binned for better visualization. As can be seen in the Figures, the contribution from each temperature component was successfully constrained by the fitting. The XMM-Newton instrumental background lines between 1.2 and 1.75 keV were not well fit with either a Gaussian, Lorentzian, or Voigt profile, so we ignored that spectral energy range in the spectral fitting. Other features present in the residual are due to the selected background model (in the case of bknpow with no gaussians), or because of cross-calibration between XMM-Newton pn, XMM-Newton MOS, and Chandra data.
@@ 4.4 The Sunyaev-Zeldovich Effect
$y_{tSZ}\equiv\int\frac{k_{\rm B}T_{\rm e}}{m_{\rm e}c^{2}}d\tau_{\rm e}=\int\frac{k_{\rm B}T_{\rm e}}{m_{\rm e}c^{2}}n_{\rm e}\sigma_{T}dl=\frac{\sigma_{T}}{m_{\rm e}c^{2}}\int P_{\rm e}dl$ (3)
@@
where $x=\frac{h\nu}{k_{B}T_{\rm{CMB}}}\approx\nu/56.8{\rm{GHz}}$ is the dimensionless frequency, $h$ is the Planck constant, $\nu$ is the observed frequency, $k_{B}$ is the Boltzmann constant, and $T_{\rm{CMB}}=2.7255{\rm{K}}$ is the CMB temperature. The $\delta_{\rm{tSZ}}(x,T_{e})$ and $\delta_{\rm{kSZ}}(x,v_{z},T_{e})$ terms correspond to the relativistic corrections for the tSZ and kSZ effects, and are dependent on the observing frequency and $T_{e}$, and in the kSZ case, also the line of sight velocity. The relativistic corrections for the tSZ are computed using the work of Itoh & Nozawa (2003), while the kSZ corrections are computed using the analytical formula in Nozawa et al. (2006). Practically, these relativistic corrections are computed from tabulated values using SZpack, whereby the corrections are on the 10% level for these 18 keV temperatures (Chluba et al., 2012, 2013).
@@ 4.5 Correction of kinetic Sunyaev-Zeldovich via ICM Physical Modeling
Figure 11: Comparison plots between the geometry-corrected X-ray pressure map (top left), the geometry-corrected trend-divided X-ray pressure map (top right), followed by Sunyaev-Zeldovich MUSTANG2 data (bottom left), next to a kinetic and relativistic SZ corrected version of the data, shown here in KelvinCMB units (bottom right). The white contours are from the radio images seen in van Weeren et al. (2017), and the green contours corresponds to the MUSTANG2 data in the bottom left. The circles represent the strong lensing subcluster centers given in Limousin et al. (2016) which were used to reproduce the kSZ+rSZ model used in Adam et al. (2017b), which are displayed in Table 8. The full FOV of the MUSTANG2 data can be seen in Appendix Sec. C.
@@
Because MUSTANG2 is dominated by the tSZ and limited to a single frequency $\sim 90\rm{GHz}$, it does not have much constraining power for the kSZ signal. As we can not constrain our own kSZ model, we instead opt to build a kSZ model for MUSTANG2. To make the correction, we use the same model parameters as ‘F2’ constrained in Adam et al. (2017b), which used the NIKA data and extra information from X-ray imaging to break degeneracies between $v_{z}$ and $\tau_{e}$. We start with the fact that previous strong lensing observations in Limousin et al. (2016) have shown that there are at least four subclusters present in MACS J0717.5+3745, listed in Table 8, and shown as circles in Fig. 11. Due to the linear dependence of $n_{e}$ in the SZ data, the 3D gas distribution is built as a sum of each independent subcluster, where each subcluster is modeled with a spherically symmetric $\beta$-model,
@@
and $n_{e0}$ is the central electron density, $r_{c}$ is the core radius, and $\beta$ is the slope of the radial profile (Cavaliere & Fusco-Femiano, 1978). Each subcluster additionally is assumed to have isothermal gas temperatures, determined using 30 arcsecond regions from XMM-Newton in Adam et al. (2017b), and constant bulk velocities, all given in Table 8. This gas distribution is then analytically integrated along the line of sight for each subcluster as,
$\tau_{e}\equiv\sigma_{T}\int_{-\infty}^{+\infty}n_{e}dl=\sqrt{\pi}\,\sigma_{T}\,n_{e0}r_{c}\frac{\Gamma\left(\frac{3}{2}\beta-\frac{1}{2}\right)}{\Gamma\left(\frac{3}{2}\beta\right)}\left[1+\left(\frac{R}{r_{c}}\right)^{2}\right]^{\frac{1}{2}-\frac{3}{2}\beta}\,$, (9)
$\frac{\Delta I_{\nu}^{\rm{model}}}{I_{0}}$ $=\sigma_{T}\sum_{i}f_{\nu}(T_{x}^{(i)})\frac{k_{B}T_{x}^{(i)}}{m_{e}c^{2}}\int n_{e}^{(i)}dl$
$+\sigma_{T}\sum_{i}g_{\nu}(T_{x}^{(i)})\frac{v_{z}^{(i)}}{c}\int n_{e}^{(i)}dl\,$, (10)
where $T_{x}$ and $v_{z}$ are scalar quantities for each subcluster.
Fig. 11 shows a comparison plot between a zoomed in view of the geometry-corrected X-ray pressure and the geometry-corrected, trend-divided pressure maps (top), followed by Sunyaev-Zeldovich MUSTANG2 data, and a kSZ and rSZ corrected version of the data. The white box corresponds to the FOV of the MUSTANG2 data, and is the same as the thermodynamical maps shown in Fig. 6 and Fig. 7, while the new smaller green box shows the zoomed-in region in the two top plots. The white contours are the radio images in van Weeren et al. (2017), and the circles represent the subcluster centers given in Table 5 which were used for the geometric correction of the maps and Table 8 used in the kSZ model.
@@
As mentioned previously, the present work does not explore pressure fluctuations (see Khatri & Gaspari, 2016; Romero et al., 2023b, 2024), but we note that deeper SZ observations with, for example, MUSTANG2, could improve upon the constraints for MACS J0717.5+3745 presented in Adam et al. (2024).